OpenStudy (anonymous):

how do I solve this? 16x^2-72x-40=0

OpenStudy (amistre64):

quadratically

OpenStudy (anonymous):

i am having problems with it I know it is quadratic

OpenStudy (amistre64):

problems in the method you chose? or in dealing with the quadratic formula?

OpenStudy (amistre64):

we can factor out a 4 4(4x^2 -18x - 10) = 0 and ignore the 4

OpenStudy (amistre64):

10*4 = 40; factors of 40 that subtract to 18 if its integerable

OpenStudy (amistre64):

1 40 2 20 ... that does it

OpenStudy (amistre64):

(x-20/4)(x+2/4) (x-5)(x+1/2) 4 (x-5) (2x+1) = 0

OpenStudy (amistre64):

im missing something i think

OpenStudy (anonymous):

no that is the answer -1/2 and 5

OpenStudy (amistre64):

ya!!

OpenStudy (amistre64):

i see what i did, I forgot to factor out another 2 along the way; but yeah, same "answers"

OpenStudy (anonymous):

2x^3-11x^2+15x=0

OpenStudy (anonymous):

how do i do this one

OpenStudy (anonymous):

there is no gcf

OpenStudy (amistre64):

there is a gcf; its "x"

OpenStudy (amistre64):

x (2x^2 -11x + 15) = 0 ; well, when x=0 for starters

OpenStudy (amistre64):

2*15 = 30 factors of 30 that add up to 11 are 6 and 5 (x-6/2) (x-5/2) x = 3 and x=-5/2 should zero it out then as well